Given, Sketch, and Solution of 

A large gun fires a shell with a muzzle velocity of 400m/sec. What is the range of the gun when it is elevated at ang angle of 30? and 50?


1
Expert's answer
2022-02-21T12:11:08-0500

Answer


Range of object projected at an angle θ

"R=\frac{v^2\sin2\theta}{g}"

At angle 30° and 50° range becomes

"R=\frac{400^2\sin2\times30°}{9.81}=14140m"


And

"R=\frac{400^2\sin2\times50°}{9.81}=16080m"